OpenStudy (ray1998):

What is the solution of the following system? 3x + 3y = 10 -9x - 9y = -20 My work so far: 3x + 3y = 10 -9x - 9y = -30 -6x -6y = -20 This problem is different from the other one, so I'm not really sure how to do it. Whoever can help me on this will receive fan and medal :) thank you all in advance!

OpenStudy (raven7266):

Are you looking for the x and y coordinates?

OpenStudy (ray1998):

Yes :)

OpenStudy (ray1998):

I don't think I'm doing this right and if I am, I'm not sure where to move on from there.

jimthompson5910 (jim_thompson5910):

@Ray1998 instead of multiplying equation one by -3, multiply it by 3 3x + 3y = 10 ----> 9x + 9y = 30 so we have this system 9x + 9y = -30 -9x - 9y = -20 add the equations to get 9x + 9y = -30 -9x - 9y = -20 -------------- 0x+0y = -50 we end up with 0 = -50 which is never true. There are no (x,y) values that make 0 = -50 true. So we have an inconsistent system that has no solutions.

jimthompson5910 (jim_thompson5910):

oops I made a typo, I put -30 instead of 30 3x + 3y = 10 ----> 9x + 9y = 30 9x + 9y = 30 -9x - 9y = -20 -------------- 0x+0y = 10 which leads to 0 = 10 the system is still inconsistent

OpenStudy (ray1998):

Okay, I think I understand what you are saying, but what happens to the 3x + 3y = 10 equation? @jim_thompson5910

OpenStudy (ray1998):

Nevermind! I see now lol. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

when I multiplied both sides by 3, it turned into 9x + 9y = 30

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (ray1998):

Thank you for all your help! This was great :)

jimthompson5910 (jim_thompson5910):

you're welcome